INTRODUCTION

This volume contains papers submitted by the participants in a workshop entitled ”Algebraic Topological Methods in Computer Science” held at Stanford University during the period July 30-August 3, 2001. The meeting was supported financially by the National Science Foundation and by the Mathematics Research Center in the Mathematics Department of Stanford University, whom we thank for their support. As is reflected in the papers, the talks at the conference were very diverse, ranging from applications of topology and category theory to concurrency in algorithms to applications in combinatorics to computational geometry. The high quality of the talks and the enthusiasm of the participants demonstrate the growing vitality of this area, and we look forward to following and participating in its development.

Gunnar Carlsson, Rick Jardine